# GeneMAN: Generalizable Single-Image 3D Human Reconstruction from Multi-Source Human Data

### v1

@@ 5.2 Quantitative Comparison
Tab. 1 and Tab. 2 provide quantitative comparisons of GeneMAN with baseline methods on CAPE [33] and in-the-wild images. As demonstrated in Tab. 2, GeneMAN outperforms all compared approaches, showcasing its superiority in generalizable human reconstruction from in-the-wild images. Notably, GeneMAN attains the highest CLIP-Similarity [42], indicating superior multi-view consistency of the reconstructed human. The significant lead in PSNR and LPIPS highlights the improved reconstruction quality of GeneMAN compared to other methods. Furthermore, GeneMAN achieves the best results in LPIPS and CLIP-Similarity, as well as comparable results in PSNR on CAPE, proving its effectiveness on laboratory dataset as well.
Figure 7: Ablation on Geometry Reconstruction and Texture Generation Stage. (a) shows the geometry initialization, (b) recovers finer details through geometry sculpting, (c) applies latent-space texturing, and (d) achieves the final reconstruction with pixel-space optimization.
Figure 8: Ablation on GeneMAN 2D and 3D Prior Models. (a) and (b) demonstrate GeneMAN 2D prior model improves multi-view consistency and texture details, while (c) and (d) illustrate enhanced pose naturalness and geometric accuracy achieved by GeneMAN 3D prior model.
Figure 9: User Study. Across all examples, $73.09\%$ of subjects prefer our method over the baselines in terms of geometry and texture, highlighting the superiority of our approach in recovering reasonable geometry and high-fidelity textures.
@@ 5.3 Qualitative Comparison
We present qualitative results on testing samples from the CAPE [33] dataset and in-the-wild images, showcased in Fig. 5 and Fig. 6, respectively. To better illustrate the reconstruction quality, we visualize the multi-view surface normals and color renderings. PIFu [46] fails to recover realistic geometry and appearance, particularly in side views. Template-based approaches, including GTA [70], TeCH [18] and SiTH [13] heavily rely on precise human pose and shape estimation (HPS). However, HPS methods [9] often produce artifacts such as “bent legs” and struggle with the shapes of children, as seen in both Fig. 5 and Fig. 6. These limitations lead to cumulative errors in geometry reconstruction that are difficult to eliminate, even with refinement. While TeCH captures more lifelike details with optimization-based refinement, the reconstructed surface appears excessively noisy, and the textures are inconsistent, especially in the back views of garments. In contrast, GeneMAN demonstrates exceptional generalizability to in-the-wild images featuring diverse clothing styles, poses and varying body proportions. Its template-free design enables superior modeling of personal belongings and loose clothing, such as basketball, skirts and dresses, while maintaining detailed, realistic geometry and natural body shapes. GeneMAN also produces high-fidelity, multi-view consistent textures for clothing and hairstyles.
@@ 5.4 User Study
To further validate the superiority of our model, we conduct a user study involving 40 participants to assess the reconstruction quality of GeneMAN compared to each baseline method across 30 test cases, evaluating both geometry and texture quality. Participants are provided with free-view rendering video for each method, and asked to select the most preferred 3D model from five randomly shuffled options. A total of $40\times 30=1200$ comparisons are conducted. As shown in Fig 9, $73.08\%$ of participants favor our method, demonstrating a significant improvement over the baselines. This indicates that our model generates more plausible and detailed geometry, along with realistic and consistent textures, confirming GeneMAN’s effectiveness in generating high-quality reconstructions.
@@ 5.5 Ablation Study
Effectiveness of Geometry Reconstruction and Texture Generation. Fig. 9 analyzes the effectiveness of each stage proposed in our framework. Compared with geometry initialization (a), geometry sculpting in (b) smooths the excessively noisy surface and recovers high-frequency details, such as clothing wrinkles and facial features. In (c), the latent space texturing yields a reasonably satisfactory result; however, the back view remains inconsistent, and the texture appears slightly blurred. To address this, (d) enhance the coarse texture into a final refined texture through pixel space texturing, optimizing the UV maps.
Effectiveness of GeneMAN 2D Prior Model. To evaluate the efficacy of our GeneMAN 2D prior model, we compare the mesh results extracted from the NeRF [36] stage utilizing the pretrained DeepFloyd-IF [45] model (“original 2D prior”) versus our GeneMAN 2D prior model (“GeneMAN 2D prior”). Note that we select the NeRF stage result for visualization, as it provides a clear view of the impact of the diffusion prior, whereas subsequent geometry and texture disentanglement may obscure this effect. As shown in Fig. 9 (a) and (b), the red-boxed regions reveal the original 2D prior results in an uneven shirt hem, with the back extending lower than the front. In contrast, GeneMAN 2D prior model ensures multi-view consistency in both geometry and texture, enhancing overall reconstruction realism.
@@ 6 Conclusion
In this paper, we present GeneMAN, a generalizable framework for single-image 3D human reconstruction. By leveraging the GeneMAN 2D and 3D prior models, trained on our multi-source human dataset, GeneMAN is capable of reconstructing high-fidelity 3D human models from in-the-wild images, accommodating varying body proportions, diverse clothing, personal belongings, and poses. Extensive experiments validate the effectiveness of our approach, demonstrating its superiority over state-of-the-art methods.
@@
This appendix serves to further enrich the discourse established in our main paper. Sec. A presents the implementation details of GeneMAN. In Sec. B, we elaborate on the details of our multi-source human dataset. Sec. C provides supplementary experimental results, including more visualizations for the qualitative comparison, discussion on geometric quantitative evaluation, and additional ablation studies conducted on our multi-source human dataset. In Sec. D, we discuss the limitations of our approach.
Appendix A Implementation Details
A.1 Camera Sampling
@@
For novel view guidance, we sample the camera distance from the range $\mathcal{U}(3.0,3.5)$. The elevation angle $\phi$ is drawn from $\mathcal{U}(-10^{\circ},45^{\circ})$, and the azimuth angle $\theta$ is uniformly sampled from $[-180^{\circ},180^{\circ})$. Additionally, the field of view (FOV) is constrained to the range $[20^{\circ},25^{\circ}]$, aligning with the fixed camera intrinsics optimized for our finetuned human diffusion models. Building upon HumanNorm [17], we segment the human body into four regions: the head, upper body, lower body, and full body. To enhance part-aware reconstruction with high-fidelity detail, we allocate a sampling probability of $0.7$ to the full body, and $0.1$ to each of the head, upper body, and lower body. We also manually adjust the camera distance for zooming in on specific parts. For instance, when refining the facial region, the camera distance is reduced to $\mathcal{U}(0.8,1.0)$. For half-body refinement (either upper or lower body), the camera distance is adjusted to $\mathcal{U}(1.5,2.0)$. Additionally, the camera center is shifted to ensure that the relevant keypoint aligns with the center of the rendered images. To achieve this, we leverage ECON [63] to identify 2D keypoints, which are then used during rasterization to back-project the coordinates into 3D space. Note that if a keypoint lies outside the image boundaries, we assign a sampling probability of zero to that point and renormalize the distribution accordingly.
A.2 Details of Each Stage
@@
In the phase of Geometry Initialization, we optimize Instant-NGP [37] from a resolution of $128$ to $384$ over the course of $5,000$ steps. The loss weights for this stage are set as follows: $\lambda_{\textnormal{r}}=1\times 10^{3},\lambda_{\textnormal{m}}=100,\lambda_{\textnormal{d}}=0.05,\lambda_{\textnormal{n}}=1,\lambda_{\textnormal{2D}}=0.1,\lambda_{\textnormal{3D}}=0.1$. Subsequently, we extract the resulting mesh from Instant-NGP with an isosurface resolution of $256$ as the geometry initialization. During the geometry sculpting stage, we refine the geometry adopting DMTet [51] at a resolution of $512$ for $10,000$ steps, enabling the capture of intricate details of humans. Inspired by HumanNorm [17], we incorporate a progressive positional encoding technique, where the mask on the position encoding for DMTet’s SDF features is gradually lifted to introduce higher-frequency components as training progresses. After $4,000$ iterations, the mask fully reveals all positions, allowing the encoding to capture both low- and high-frequency details. Empirically, we observe that omitting progressive positional encoding results in noisy surface reconstructions. The loss weights are set as follows: $\lambda_{r}=5\times 10^{3},\lambda_{vgg}=1\times 10^{3},\lambda_{sdf}=1.5\times 10^{3}$ and $\lambda_{sds}=1.0$. To ensure consistency between the rendered front view and the input image, we apply a relatively small guidance scale of $20$ for novel view generation using the normal- and depth-adapted diffusion models [17]. The noise timestep $t$ is sampled from $\mathcal{U}(0.02,0.8)$. Besides, we adopt the AdamW [31] optimizer with a base learning rate of $0.01$ during Geometry Initialization and $2\times 10^{-5}$ during Geometry Sculpting.
Multi-Space Texture Refinement. For latent space texture refinement, we employ SDS optimization [41] in the latent space to optimize the coarse texture at an image resolution of $1024\times 1024$ for $12,000$ steps. The loss weights for the coarse texture stage are configured as follows: $\lambda^{color}_{rgb}=1\times 10^{3},\lambda^{color}_{m}=100,\lambda^{color}_{2D}=0.1,\lambda^{color}_{3D}=0.1$. Following this, we perform pixel-space texture refinement by optimizing the UV texture map for an additional 1,000 steps to achieve finer texture details. To ensure that the added noise does not corrupt the original content while retaining the capacity to enhance image details, we empirically set the starting timestep $t_{\textnormal{start}}$ to $0.05$ in Eq.8 in the main text. The weight for the VGG loss [53] in the pixel space texture refinement is set to $\lambda_{LP}=0.01$. We adopt the AdamW [31] optimizer with a base learning rate of $0.01$ in the coarse texture stage and $0.02$ in the fine texture stage.
Appendix B Details of Multi-Source Human Dataset
@@
Figure 10: Image Cropping Synthesis. We perform half-body and three-quarters-length cropping on the multi-view renderings of 3D scans to generate synthetic cropped images.
Appendix C Additional Experiment Results
C.1 More Qualitative Results
For a more comprehensive geometric evaluation, we incorporated four additional SOTA human geometry reconstruction methods for comparison: PaMIR [72], ICON [62], ECON [63], and SIFU [71]. The geometric comparison results are presented in Fig. 13 and Fig. 14. It can be seen that our model effectively recovers natural human poses, accommodates loose clothing, and excels at reconstructing images with diverse body ratios. In addition, we provide more qualitative comparisons with state-of-the-art methods: PIFu [46], GTA [70], TeCH [18], SiTH [13] as shown in Fig. 15 to Fig. 21. Our method surpasses the compared methods in generating consistent, highly realistic textures with exceptional fidelity.
C.2 Quantitative Geometric Evaluation
We also perform a quantitative geometric comparison with baseline methods [46, 72, 62, 63, 70, 18, 13, 71] on the CAPE [33] dataset to assess geometry reconstruction quality. Specifically, we report two commonly used metrics: Chamfer distance (bi-directional point-to-surface) and Point-to-Surface (P2S) distance (1-directional point-to-surface), both measured in centimeters, between the ground truth scans and the reconstructed meshes. Additionally, to assess the fidelity of reconstructed local details, we calculate the $\mathcal{L}_{2}$ error between normal images rendered from the reconstructed and GT surfaces. These renderings are obtained by rotating the camera around the surfaces at angles of $\{0^{\circ},120^{\circ},240^{\circ}\}$ relative to the frontal view.

### v2

@@ 5.2 Quantitative Comparison
Tab. 1 provides quantitative metrics of GeneMAN with baseline methods on in-the-wild images and CAPE [33]. For in-the-wild images, GeneMAN outperforms all competing approaches, showcasing its superiority in generalizable human reconstruction, with the highest CLIP-Similarity [43] indicating better multi-view consistency. It also leads in PSNR and LPIPS, highlighting its improved reconstruction quality. On CAPE, GeneMAN continues to excel, achieving the best results in LPIPS and CLIP-Similarity while maintaining competitive performance in PSNR, proving its effectiveness on laboratory dataset as well.
Figure 8: Ablation on Geometry Reconstruction and Texture Generation. (a) shows the geometry initialization, (b) recovers finer details through geometry sculpting, (c) applies latent-space texturing, and (d) achieves the final reconstruction with pixel-space optimization.
Figure 9: Ablation on GeneMAN 2D and 3D Prior Models. (a) and (b) demonstrate the 2D prior model improves multi-view consistency and texture details, while (c) and (d) illustrate enhanced pose naturalness and geometric accuracy achieved by the 3D prior model.
Figure 10: User Study. Across all examples, $73.09\%$ of subjects prefer our method over the baselines in terms of geometry and texture, highlighting the significant superiority of our approach in recovering accurate geometry and high-fidelity appearances.
@@ 5.3 Qualitative Comparison
We present qualitative results on testing samples from in-the-wild images and the CAPE [33] dataset, showcased in Fig. 6 and Fig. 7, respectively. To better illustrate the reconstruction quality, we visualize the multi-view surface normals and color renderings. PIFu [47] fails to recover realistic geometry and appearance, particularly in side views. Template-based approaches, including GTA [71], TeCH [18] and SiTH [13] heavily rely on precise human pose and shape estimation (HPS). However, HPS methods [9] often produce artifacts such as “bent legs” and struggle with the shapes of children, as seen in both Fig. 6 and Fig. 7. These limitations lead to cumulative errors in geometry reconstruction that are difficult to eliminate, even with refinement. While TeCH captures more lifelike details with optimization-based refinement, the reconstructed surface appears excessively noisy, and the textures are inconsistent, especially in the back views of garments. In contrast, GeneMAN demonstrates exceptional generalizability to in-the-wild images featuring diverse clothing styles, poses and varying body proportions. Its template-free design enables superior modeling of personal belongings and loose clothing, such as basketball, skirts and dresses, while maintaining detailed, realistic geometry and natural body shapes. GeneMAN also produces high-fidelity, multi-view consistent textures for clothing and hairstyles. In Fig. 5, we present additional results to highlight the robust reconstruction capabilities of GeneMAN under complex poses. More results on complex poses are provided in the supplementary.
@@ 5.4 User Study
We conduct a user study involving 40 participants to assess the reconstruction quality of GeneMAN compared to each baseline method across 30 test cases, evaluating both geometry and texture quality. Participants are provided with free-view rendering video for each method, and asked to select the most preferred 3D model from five randomly shuffled options. A total of $40\times 30=1200$ comparisons are conducted. As shown in Fig. 10, $73.08\%$ of participants favor our method, demonstrating a significant improvement over the baselines. This indicates that our model generates more plausible and detailed geometry, along with realistic and consistent textures, confirming GeneMAN’s effectiveness in generating high-quality reconstructions.
@@ 5.5 Ablation Study
Effectiveness of Geometry Reconstruction and Texture Generation. Fig. 10 analyzes the effectiveness of each stage proposed in our framework. Compared with geometry initialization (a), geometry sculpting in (b) smooths the excessively noisy surface and recovers high-frequency details, such as clothing wrinkles and facial features. In (c), the latent space texturing yields a reasonably satisfactory result; however, the back view remains inconsistent, and the texture appears slightly blurred. To address this, (d) enhance the coarse texture into a final refined texture through pixel space texturing, optimizing the UV map.
Effectiveness of GeneMAN 2D Prior Model. To evaluate the efficacy of our GeneMAN 2D prior model, we compare the mesh results extracted from the NeRF [36] stage utilizing the pretrained DeepFloyd-IF [46] model (“original 2D prior”) versus our GeneMAN 2D prior model (“GeneMAN 2D prior”). Note that we select the NeRF stage output for visualization to clearly demonstrate the impact of the diffusion prior, whereas subsequent geometry and texture disentanglement may obscure this effect. As illustrated in Fig. 10 (a) and (b), the red-boxed regions reveal that the original 2D prior results in an uneven shirt hem, with the back extending lower than the front. In contrast, GeneMAN 2D prior model maintains multi-view consistency in both geometry and texture, enhancing overall reconstruction.
@@ 6 Conclusion
In this paper, we present GeneMAN, a generalizable framework for single-image 3D human reconstruction. By leveraging the GeneMAN 2D and 3D prior models, trained on a large-scale, multi-source human dataset, GeneMAN is capable of reconstructing high-fidelity 3D human models from in-the-wild images, accommodating varying body proportions, diverse clothing, personal belongings, and complex poses. Extensive experiments validate the effectiveness of our approach, demonstrating its superiority over state-of-the-art methods.
[…]
@@
This appendix serves to further enrich the discourse established in our main paper. Sec. A offers additional information to supplement the related work. Sec. B presents the implementation details of GeneMAN. In Sec. C, we elaborate on the details of our multi-source human dataset. Sect. D provides supplementary experimental results, including more visualizations for the qualitative comparison, additional results of GeneMAN on in-the-wild images and CAPE [33], discussion on geometric quantitative evaluation, and ablation studies conducted on our multi-source human dataset. In Sec. E, we discuss the limitations of our approach.
Appendix A Supplement to Related Work
A.1 Text-to-3D Generation
The rapid advancements in 2D image generation have also significantly accelerated progress in the field of text-to-3D generation. DreamFusion [41] and Magic3D [25] have demonstrated that utilizing pretrained 2D text-to-image diffusion models [45, 44] as guidance can greatly enhance the optimization of 3D representations through score distillation sampling (SDS). Trained on large-scale 2D image datasets, these text-to-image models possess excellent capabilities in providing comprehensive 2D knowledge and hallucinating unseen scenes based on specific prompts. Nevertheless, as these models are limited to 2D knowledge, the aforementioned text-to-3D methods are susceptible to multi-view consistency issues. MVDream [53] improves 3D consistency by training a multi-view diffusion model that offers a robust multi-view prior. However, these methods are not well-suited for direct image-to-3D reconstruction tasks because images cannot be accurately captured through textual descriptions, leading to inconsistencies in color and texture across the generated assets.
A.2 Singe Image-to-3D Reconstruction
Single image-to-3D reconstruction has greatly benefited from the thriving advancement of diffusion models. Zero-1-to-3 [28] trains a view-dependent diffusion model on Objaverse [8] by explicitly incorporating the camera parameters into the diffusion process. It enables zero-shot image-conditioned novel view synthesis and facilitates general image-to-3D reconstruction with consistent geometric view. Magic123 [42] and DreamCraft3D [58] integrate the 3D prior from Zero-1-to-3 [28] with 2D prior from text-to-image diffusion model, employing a coarse-to-fine two-stage optimization to achieve high-quality 3D reconstruction. LRM [16] leverages both synthetic data from Objaverse [8] and multi-view data from MVImageNet [67] to train a transformer-based 3D architecture, enhancing its generalizability for 3D reconstruction. However, these general image-to-3D approaches often yield poor results in human reconstruction, resulting in inaccurate geometry of human bodies and the neglect of intricate details, such as facial features and clothing. This limitation arises from the lack of human-specific priors.
Appendix B Implementation Details
B.1 Camera Sampling
@@
For novel view guidance, we sample the camera distance from the range $\mathcal{U}(3.0,3.5)$. The elevation angle $\phi$ is drawn from $\mathcal{U}(-10^{\circ},45^{\circ})$, and the azimuth angle $\theta$ is uniformly sampled from $[-180^{\circ},180^{\circ})$. Additionally, the field of view (FOV) is constrained to the range $[20^{\circ},25^{\circ}]$, aligning with the fixed camera intrinsics optimized for our finetuned human diffusion models. Building upon HumanNorm [17], we segment the human body into four regions: the head, upper body, lower body, and full body. To enhance part-aware reconstruction with high-fidelity detail, we allocate a sampling probability of $0.7$ to the full body, and $0.1$ to each of the head, upper body, and lower body. We also manually adjust the camera distance for zooming in on specific parts. For instance, when refining the facial region, the camera distance is reduced to $\mathcal{U}(0.8,1.0)$. For half-body refinement (either upper or lower body), the camera distance is adjusted to $\mathcal{U}(1.5,2.0)$. Additionally, the camera center is shifted to ensure that the relevant keypoint aligns with the center of the rendered images. To achieve this, we use an off-the-shelf tool to identify 2D keypoints, which are then used during rasterization to back-project the coordinates into 3D space. Note that if a keypoint lies outside the image boundaries, we assign a sampling probability of zero to that point and renormalize the distribution accordingly.
B.2 Details of Each Stage
@@
In the phase of Geometry Initialization, we optimize Instant-NGP [37] from a resolution of $128$ to $384$ over the course of $5,000$ steps. The loss weights for this stage are set as follows: $\lambda_{\textnormal{r}}=1\times 10^{3},\lambda_{\textnormal{m}}=100,\lambda_{\textnormal{d}}=0.05,\lambda_{\textnormal{n}}=1,\lambda_{\textnormal{2D}}=0.1,\lambda_{\textnormal{3D}}=0.1$. Subsequently, we extract the resulting mesh from Instant-NGP with an isosurface resolution of $256$ as the geometry initialization. During the geometry sculpting stage, we refine the geometry adopting DMTet [52] at a resolution of $512$ for $3,000$ steps, enabling the capture of intricate details of humans. Inspired by HumanNorm [17], we incorporate a progressive positional encoding technique, where the mask on the position encoding for DMTet’s SDF features is gradually lifted to introduce higher-frequency components as training progresses. After $2,000$ iterations, the mask fully reveals all positions, allowing the encoding to capture both low- and high-frequency details. Empirically, we observe that omitting progressive positional encoding results in noisy surface reconstructions. The loss weights are set as follows: $\lambda_{r}=5\times 10^{3},\lambda_{vgg}=1\times 10^{3},\lambda_{sdf}=1.5\times 10^{3}$ and $\lambda_{sds}=1.0$. To ensure consistency between the rendered front view and the input image, we apply a relatively small guidance scale of $20$ for novel view generation using the normal- and depth-adapted diffusion models [17]. The noise timestep $t$ is sampled from $\mathcal{U}(0.02,0.8)$. Besides, we adopt the AdamW [31] optimizer with a base learning rate of $0.01$ during Geometry Initialization and $2\times 10^{-5}$ during Geometry Sculpting.
Multi-Space Texture Refinement. For latent space texture refinement, we employ SDS optimization [41] in the latent space to optimize the coarse texture at an image resolution of $1024\times 1024$ for $10,000$ steps. The loss weights for the coarse texture stage are configured as follows: $\lambda^{color}_{rgb}=1\times 10^{3},\lambda^{color}_{m}=100,\lambda^{color}_{2D}=0.1,\lambda^{color}_{3D}=0.1$. Following this, we perform pixel-space texture refinement by optimizing the UV texture map for an additional 1,000 steps to achieve finer texture details. To ensure that the added noise does not corrupt the original content while retaining the capacity to enhance image details, we empirically set the starting timestep $t_{\textnormal{start}}$ to $0.05$ in Eq.8 in the main text. The weight for the VGG loss [54] in the pixel space texture refinement is set to $\lambda_{LP}=0.01$. We adopt the AdamW [31] optimizer with a base learning rate of $0.01$ in the coarse texture stage and $0.02$ in the fine texture stage.
Appendix C Details of Multi-Source Human Dataset
Figure 11: Image Cropping Synthesis. We apply half-body and three-quarters-length cropping to the multi-view renderings of 3D scans to generate synthetic cropped images.
@@ Appendix C Details of Multi-Source Human Dataset
Appendix D Additional Experiment Results
D.1 More Qualitative Results
For a more comprehensive geometric evaluation, we incorporated six additional SOTA human geometry reconstruction methods for comparison: PIFuHD [48], PaMIR [73], ICON [63], ECON [64], PHORHUM [4], and SIFU [72]. The geometric comparison results are presented in Fig. 17 and Fig. 18. It can be seen that our model effectively recovers natural human poses, accommodates loose clothing, and excels at reconstructing images with diverse body ratios. In addition, we provide more qualitative comparisons with state-of-the-art methods: PIFu [47], GTA [71], TeCH [18], SiTH [13] as shown in Fig. 19 to Fig. 25. Our method surpasses the compared methods in generating consistent, highly realistic textures with exceptional fidelity.
Besides, in Fig. 13, we conduct additional comparisons with two generalizable methods: HumanLRM [61] and IDOL [74], both of which are feed-forward models trained on large-scale human datasets. As the code of Human-LRM is unavailable, we use the examples provided in its paper. Our model authentically reconstructs 3D humans that closely align with the front view, while IDOL fails to preserve facial details. Our model also produces natural poses with reasonable geometry (see the legs of the first and the third subjects in Fig. 13). Moreover, it achieves high-fidelity appearance with view-consistent details, significantly outperforming HumanLRM.
We provide additional visualization results, with a particular focus on complex poses, in Fig. 15 (in-the-wild images) and Fig. 16 (CAPE [33]). These results highlight the strong generalization ability of GeneMAN to challenging poses.
Figure 13: Additional Comparison on in-the-wild Cases with HumanLRM [61] and IDOL [74]. We conduct experiments on the examples reported by HumanLRM. Best view with zoomed in.
D.2 Quantitative Geometric Evaluation
Following prior work, we perform a quantitative geometric comparison with baseline methods [47, 73, 63, 64, 71, 18, 13, 72] on the CAPE [33] dataset to assess geometry reconstruction quality. Specifically, we report two commonly used metrics: Chamfer distance (bi-directional point-to-surface) and Point-to-Surface (P2S) distance (1-directional point-to-surface), both measured in centimeters, between the ground truth scans and the reconstructed meshes. Additionally, to assess the fidelity of reconstructed local details, we calculate the $\mathcal{L}_{2}$ error between normal images rendered from the reconstructed and GT surfaces, referred to as Normal Consistency (NC). These renderings are obtained by rotating the camera around the surfaces at angles of $\{0^{\circ},120^{\circ},240^{\circ}\}$ relative to the frontal view.